OpenStudy (anonymous):

Write the equation of the line in point-slope form for the line through the given point (4, -6) and with m = 3/5 a. y + 6 = 3/5x - 4 b. y - 6 = 3/5 (x+4) c. y + 6 = 3/5 (x-4) d. y - 4 = 3/5 (x+6)

OpenStudy (anonymous):

y -y1 = m (x-x1) where m is slope and x1, y1 is a point on the line y - (-6) = 3/5 (x- 4) y + 6 = 3/5 (x-4) answer is c